# Agent-based modeling for realistic reproduction of human mobility and contact behavior to evaluate test and isolation strategies in epidemic infectious disease spread

### v1

@@ Abstract
{highlights}
An agent-based model including realistic individual mobility and transmission of infections
Efficient and scalable implementation for simulations of several million agents
Our results suggest that symptom-independent testing strategies have limited effect if symptomatic control is low.
Our results suggest that with sufficient symptomatic control, even short quarantines have a substantial effect on the mitigation of disease dynamics
@@ 1 Introduction
Mathematical models are invaluable assets in many application domains, as they allow to proactively study alternative scenarios of future developments. For infectious disease dynamics, models based on simple systems of ordinary differential equations (ODEs) are popular tools that allow a quick and easy assessment of situations at hand; see, e.g., several chapters in [3]. However, these models are often overly simplified and use several unrealistic assumptions. While spatial or demographic stratification can already be resolved within larger ODE systems [4] or recent semi-discrete approaches [5, 6], unrealistic distribution times [7, 8] can be overcome by using a linear chain trick [9] or its generalization [10] or even Integro-differential equation-based systems [11]. Nevertheless, all these models remain aggregated models and only allow limited heterogeneity in individual reaction to transmission or infection. On the other hand, contact-network, agent- (ABMs) or individual-based models [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] allow using arbitrary levels of heterogeneity in individual human behavior and reaction to infection. Note that [28] makes a distinction between agent-based models (with movements and interactions of individuals) and contact-network models (with a detailed network of social interactions between individuals) which is not made here.
@@
The authors of [12] presented ground-breaking and promising ABMs on HPC infrastructure applied to, e.g., rumor spreading. The authors of [13] presented another very interesting multicore ABM and shed light on many different and costly aspects in agent-based modeling. From the geolocations in a 24h interval, the authors of [14] derived a social contact network with edge weights given by the duration of contacts. In 2018, the powerful platform EMOD, allowing for environmental, vector, airborne or sexual transmittable diseases, was presented by [15]. In [16], innovation spreading was modeled through an ABM. In [17], the authors modified a previously developed simulation tool, originally designed for Influenza studies, to analyze COVID-19 in the UK and the US. Their model includes detailed interactions within households, workplaces, schools, and the wider community, forming a structured contact network that reflects the primary contexts for disease transmission. The level of adherence to interventions, such as case isolation and quarantine, was incorporated into the model with specific compliance rates, influencing the simulation outcomes and reflecting realistic variations in public response. The model of [24] was developed using ChiSIM [25], a versatile community model initially designed for exploring a range of social dynamics in Chicago, including the transmission of infectious diseases. This model utilizes the Repast ABM framework [12] to facilitate a distributed memory simulation that operates on high-performance computing (HPC) systems. In each simulated hour, agents choose a single activity from their predefined activity schedules to be performed at a specific location. Additionally, agents have the option to alter their behavior and wear masks, practice social distancing, or opting to stay home, which serves as a surrogate for quarantine measures. The calibration of the model, particularly the agents’ stay-at-home behavior and the disease model parameters, was aligned with data on hospital bed usage and deaths attributed to COVID-19. The COVID-19 Modeling Kit (COMOKIT) [26] is an agent-based modeling software on the GAMA platform [27], designed to analyze COVID-19 response strategies in various contexts. COMOKIT combines models of person-to-person and environmental transmission, a model of individual epidemiological status evolution, an agenda-based one hour time step model of human mobility, and an intervention model. It integrates models for transmission, policy, and individual behavior, and is adaptable for different scenarios using minimal initial data. Despite its effectiveness, the current version faces challenges in scaling up for larger populations or fully representing social dynamics in outdoor group activities. While this list is nonexhaustive, many other authors used ABMs to simulate COVID-19, see, e.g., [18, 19, 20, 21, 22, 23, 29].
@@
The model presented in this study is a trip-based ABM, which means that it explicitly realizes mobility and allows microscopic contact patterns inside locations. Through this approach, NPIs can be realized in a fully natural way by enforcing venues to be closed, capacity-restricted, or by requiring vaccinations or negative tests of any entering agents. In contrast to equation- or mean-field models, it also allows for the study of heterogeneous individual immunity and transmission dynamics and enables settings like face masks, testing, isolation, and compliance. On one hand, our ABM allows for efficient computations of city-scale infectious disease dynamics on a single laptop. On the other hand, our model has been parallelized, both, on an intra- and inter-simulation level such that large-scale simulations can be efficiently conducted on a HPC infrastructure.
@@ 7      $t\leftarrow t+\Delta t$
Figure 1: Two exemplary time steps of the ABM simulation with five agents. From top left to bottom right, persons interact and potentially infect each other at a location. Then, agents potentially move to other location types, e.g., from a supermarket to work, or if they have severe symptoms. After a time step, the infection state can progress and potentially severely infected agents go to the hospital.
@@ 2.2 Locations
    a reduction factor for the likelihood of entering the location (when restrictions are in place) and an information on location-specific restrictions such as the maximum number of persons allowed to enter the location,
@@ 2.3.1 Mobility rules
Our current implementation on mobility rules is rather basic. Depending on the application, customs and cultural or regional practices like Easter holidays, Christmas or particular New Year’s events, more advanced schemes have to be provided by the user.
@@ 2.4 Contacts
The mobility and presence of agents at locations as described in the previous sections define sets of potential contacts of the agents. One option to realize close contacts, is the use of human locomotion models inside locations, see, e.g., [39, 40]. In general, the average number of daily contacts individuals have inside an ABM should be in line with aggregated numbers as obtained by [41, 42, 43] and used in ODE-based models. However, as these numbers are reported on an aggregated level, it is impossible to obtain a unique disaggregated, individual number of contacts per location. Furthermore, a trip-based ABM will often yield particular situations in which the average number of contacts cannot be reached by individual agents. On the other hand, individual-based data sets cannot be obtained due to data protection regulations.
@@ 2.5.1 Viral load and infection states
We consider two main representatives of an infection, viral load and (symptomatic) infection states. The viral load time course of an infected agent $p$ describes the amount of virus particles carried by the host. It is modeled by a continuous function in time. As observed in [44], viral load time courses consist of a rapid exponential increase of virus particles until a peak level, followed by a shallow exponential decrease until the virus is cleared by the immune system. Mathematically, we model the time of peak viral load as $t_{\textrm{P,p}}=t_{\textrm{T,p}}+\frac{v_{\textrm{P,p}}}{v_{\textrm{I,p}}}$, where $t_{\textrm{T,p}}$ is the time of transmission and $v_{\textrm{I,p}},v_{\textrm{P,p}}>0$ are the viral load increase rate and peak, respectively. The logarithmic viral load $v_{p}$ can be described as
@@
Figure 2: Viral load (left) and different, corresponding viral sheds (right). The presented viral sheds are given for varying factor $s_{\textrm{f,p}}$ in Eq. (2), ranging from 0.01 (blue) to 0.28 (red).
@@
While an early systematic review [45] and, e.g., [46] found higher SARS-CoV-2 viral loads associated with more severe COVID-19, a more recent systematic review [47] was inconclusive and found contradictory studies on the relationship between viral load and COVID‐19 severity. Accordingly, we model symptoms independently of the viral load. However, the time of symptom onset was observed to be close to the time of peak viral load [48, 49]. hen setting parameters, this is taken into account.
@@
Figure 3: Infection state transition model. Rounded gray borders indicate noninfectious states while red angular borders indicate symptom states that are used for infectious agents only. Orange to violet colors indicate the symptomatic severity and dashed borders indicates if a person is considered as symptomatic (for testing schemes).
@@ 2.5.2 Transmissions
We consider the viral shed as the proxy of the infectiousness of an agent. This approach is advantageous because viral shedding directly reflects the amount of virus being released by an individual, which has been shown to have a large impact on transmission potential [50]. Unlike symptom-based measures, which may lag behind or be absent in asymptomatic cases, viral shedding captures infectiousness even before symptoms appear or in cases where individuals remain asymptomatic throughout the infection [51, 48]. This proxy also allows for a more granular and dynamic assessment of infectiousness over time. As studies [48, 49] have observed that viral load peaks close to symptom onset, symptom states alone would underestimate the transmission risk during the early, highly infectious phase [44, 52].
@@
where $s_{\textrm{f,p}}$ is the personal viral shed factor for the agent $p$. The viral shed factor is drawn randomly from a probability distribution for each agent and can vary greatly, c.f. [52, Fig. 3]; also see Fig. 2. Similarly to the viral load above, we neglect any viral shed outside of the time interval $[t_{\textrm{T,p}},t_{\textrm{C,p}}]$. Furthermore, we consider only agents which are not Susceptible, Exposed, Recovered and Dead to be infectious. For instance, this means that potential viral loads larger zero are cut on symptomatic recovery. As this leads to a potential cutoff of $s_{\textrm{p}}(t)$, this needs to be taken into account, when determining $\alpha$ and $\beta$. In our simulations, cutting viral loads for recovered people happens mostly when the viral shed is already low, thus leading to a minimal loss of total infectiousness.
@@ 10             Draw new course of infection
Agent A’s exposure rate is further adjusted by seasonality and NPIs, resulting in a final exposure rate of $0.128$ for Agent A. Given that Agent A is wearing a mask that provides $m_{\textrm{p}}=85$ protection ($0.151-m_{\textrm{p}}(t)=$ 0.15), the individual infection rate for Agent A is calculated as $0.038$, as detailed in Table 2. This infection rate will be used to determine the likelihood that Agent A contracts the virus in this specific time step.
@@ 2.6.1 Testing and Isolation
Figure 4: An exemplary TestingStrategy with three different TestingSchemes and different TestingCriteria (bottom part, light blue).
@@
Figure 5: Example of TestingSchemes 1 and 2 from Fig. 4. In TestingScheme 1, agents are obliged to test themselves before going to work. It is sufficient to use a Rapid-Antigen Test if they do not show symptoms. All agents who are not symptomatic, i.e., the first three agents in the bottom row, do an antigen test from which the first is false-negative, the second positive and the third negative. Through TestingScheme2, a PCR-Test is required if the agent shows symptoms. The last remaining, symptomatic agent in the bottom row performs a PCR test which turns out the be positive. The two positive tests result in quarantining at home.
@@ 2.6.2 Venue restrictions and closures
First, the model allows to completely close (particular) locations. Second, it is possible to adjust the capacity of locations, and third, to adjust the likelihood of particular agents to move to a location. By this, we can model a reduced percentage of persons to enter such as only 50 % of pupils attending school on a particular day. If a capacity threshold is reached, further agents are not allowed to enter the location. For every location, the particular contact matrix can be scaled to simulate contact reduction and social distancing.
@@ 2.6.3 Masks and Compliance
Masks are a crucial NPI to reduce the spread of the virus but their overall effectiveness is highly dependent on correct wearing and compliance [59]. Therefore, our model also implements the usage of different types of Masks such that locations can require agents to use a particular mask – in this case stricter or more effective masks can also be worn to enter a location. In our simulations, a mask reduces the amount of virus particles emitted from an infectious agent and, on the other hand, it reduces the amount of virus particles received by other agents. For details, see Eqs. (3) and (5). Through this implementation on the individual level, the effectiveness of mask wearing by highly vulnerable persons could be studied with the software. We, furthermore, introduced individual compliance to mask wearing. With the maximum negative compliance value, a person never adheres to any mask mandate, while persons with maximum positive compliance wear a mask even if it is not mandatory. Furthermore, while compliance can be agent-specific, one agent can have different compliance values for wearing masks in different location types.

### v2

@@ 1 Introduction
Mathematical models are invaluable assets in many application domains, as they allow to proactively study alternative scenarios of future developments. For infectious disease dynamics, models based on simple systems of ordinary differential equations (ODEs) are popular tools that allow a quick and easy assessment of situations at hand; see, e.g., several sections in [3]. However, these models are often overly simplified and use several unrealistic assumptions. While spatial or demographic stratification can already be resolved within larger ODE systems [4] or recent semi-discrete approaches [5, 6], unrealistic distribution times [7, 8] can be overcome by using a linear chain trick [9] or its generalization [10] or even Integro-differential equation-based systems [11]. Nevertheless, all these models remain aggregated models and only allow limited heterogeneity in individual reaction to transmission or infection. On the other hand, contact-network, agent- (ABMs) or individual-based models allow using arbitrary levels of heterogeneity in individual human behavior and reaction to infection.
Very early works laying the foundations to modern agent-based modeling can be traced back to the 1940s to 1950s with John von Neumann’s work on self-reproducing automata [12]. As computational power increased during the late 20th century, ABMs gained more prominence in various fields, such as in social sciences with Thomas Schelling’s segregation model [13] showing the potential of ABMs in the 1970s. As computational capabilities allowed more sophisticated models later on, various ABMs were developed in other domains allowing researchers to explore complex market dynamics, population dynamics, or social phenomena; see [14, 15, 16]. Over the past few decades, increasingly detailed data on individual behaviors and heterogeneous contact patterns have become available. This advancement has enabled the construction of synthetic populations and contact networks, which serve as the foundation for not only our agent-based models but also for most agent-based approaches in infectious disease analysis [17, 18, 19, 20, 21, 22]. Finally, with the advent of modern computers and supercomputers, highly detailed and parametrized ABMs became feasible to be applied, in particular, to infectious disease spread of different pathogens; see, e.g.,  [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] for different developments of the last few decades and recent novel research emerging from the SARS-CoV-2 pandemic. Note that [39] makes a distinction between agent-based models (with movements and interactions of individuals) and contact-network models (with a detailed network of social interactions between individuals) which is not made here.
@@
The authors of [23] presented ground-breaking and promising ABMs on HPC infrastructure applied to, e.g., rumor spreading. The authors of [24] presented another very interesting multicore ABM and shed light on many different and costly aspects in agent-based modeling. From the geolocations in a 24h interval, the authors of [25] derived a social contact network with edge weights given by the duration of contacts. In 2018, the powerful platform EMOD, allowing for environmental, vector, airborne or sexual transmittable diseases, was presented by [26]. In [27], innovation spreading was modeled through an ABM. In [28], the authors modified a previously developed simulation tool, originally designed for Influenza studies, to analyze COVID-19 in the UK and the US. Their model includes detailed interactions within households, workplaces, schools, and the wider community, forming a structured contact network that reflects the primary contexts for disease transmission. The level of adherence to interventions, such as case isolation and quarantine, was incorporated into the model with specific compliance rates, influencing the simulation outcomes and reflecting realistic variations in public response. The model of [35] was developed using ChiSIM [36], a versatile community model initially designed for exploring a range of social dynamics in Chicago, including the transmission of infectious diseases. This model utilizes the Repast ABM framework [23] to facilitate a distributed memory simulation that operates on high-performance computing (HPC) systems. In each simulated hour, agents choose a single activity from their predefined activity schedules to be performed at a specific location. Additionally, agents have the option to alter their behavior and wear masks, practice social distancing, or opting to stay home, which serves as a surrogate for quarantine measures. The calibration of the model, particularly the agents’ stay-at-home behavior and the disease model parameters, was aligned with data on hospital bed usage and deaths attributed to COVID-19. The COVID-19 Modeling Kit (COMOKIT) [37] is an agent-based modeling software on the GAMA platform [38], designed to analyze COVID-19 response strategies in various contexts. COMOKIT combines models of person-to-person and environmental transmission, a model of individual epidemiological status evolution, an agenda-based one-hour time step model of human mobility, and an intervention model. It integrates models for transmission, policy, and individual behavior, and is adaptable for different scenarios using minimal initial data. Despite its effectiveness, the current version faces challenges in scaling up for larger populations or fully representing social dynamics in outdoor group activities. While this list is nonexhaustive, many other authors used ABMs to simulate COVID-19, see, e.g., [29, 30, 31, 32, 33, 34, 40].
@@
The model presented in this study is a trip-based ABM, which means that it explicitly realizes mobility and allows microscopic contact patterns inside locations. Through this approach, NPIs can be realized in a fully natural way by enforcing venues to be closed, capacity-restricted, or by requiring vaccinations or negative tests of any entering agents. In contrast to equation- or mean-field models, it also allows for the study of heterogeneous individual immunity and transmission dynamics and enables settings like face masks, testing, isolation, and compliance. On one hand, our ABM allows for efficient computations of city-scale infectious disease dynamics on a single laptop. On the other hand, our model has been parallelized, both, on an intra- and inter-simulation level such that large-scale simulations can be efficiently conducted on an HPC infrastructure.
@@ 2 Material and methods
Furthermore, many technical details and formulas are presented in Section 2.5.2, Section 2.6.1, Section 2.8 and Section 2.9. Although these details are important to fully understand the entire model and its capabilities, a general understanding can be obtained without the technical details, in particular through the support of Algorithm 1 to Algorithm 3 and Fig. 1 to Fig. 4.
@@ 7      $t\leftarrow t+\Delta t$
Figure 1: Two exemplary time steps of the ABM simulation with five agents. From top left to bottom right, persons interact and potentially infect each other at a location. Then, agents potentially move to other location types, e.g., from a supermarket to work, or to the hospital if they have severe symptoms. After a time step, the infection state can progress, e.g., from mildly infected to severely infected.
@@ 2.2 Locations
    a reduction factor for the likelihood of entering the location (when restrictions are in place) and information on location-specific restrictions such as the maximum number of persons allowed to enter the location,
@@ 2.3.1 Mobility rules
Our current implementation of mobility rules is rather basic. Depending on the application, customs and cultural or regional practices like Easter holidays, Christmas or particular New Year’s events, more advanced schemes have to be provided by the user.
@@ 2.4 Contacts
The mobility and presence of agents at locations as described in the previous sections define sets of potential contacts of the agents. One option to model close contacts, is the use of human locomotion models inside locations, see, e.g., [50, 51]. In general, the average number of daily contacts individuals have inside an ABM should be in line with aggregated numbers as obtained by [52, 53, 54] and used in ODE-based models. However, as these numbers are reported on an aggregated level, it is impossible to obtain a unique disaggregated, individual number of contacts per location. Furthermore, a trip-based ABM will often yield particular situations in which the average number of contacts cannot be reached by individual agents. On the other hand, individual-based data sets cannot be obtained due to data protection regulations.
@@ 2.5 Transmission and infection model
As this section is of high technical detail, we provide a quick summary beforehand which enables a quick, shallow understanding of our infection and transmission model: We describe each agent’s infectiousness time-dependent, beginning at zero at infection transmission, rising during the infection until the peak level of viral load, and then slowly dropping again to zero until the virus is cleared from the host. During a contact with an infected agent, this time-dependent infectiousness is evaluated, and the chance of infection is derived, taking into account further properties of both the transmitting and receiving agent, such as the use of masks and an individual protection level from prior infections or vaccinations. In addition, each agent undergoes a series of symptomatic states, beginning at the time of positive virus transmission, and ending in a recovered or dead state. This state directly influences whether and when the agent visits the hospital or is tested. Building on this quick overview, we now explain the realization of the infection and transmission model in detail, starting with the twofold representation of an infection through viral load and infection states in Section 2.5.1 and how agents infect each other in Section 2.5.2.
@@ 2.5.1 Viral load and infection states
We consider two main representatives of an infection, viral load and (symptomatic) infection states. The viral load time course of an infected agent $p$ describes the amount of virus particles carried by the host. It is modeled by a continuous function in time. As observed in [55], viral load time courses consist of a rapid exponential increase in virus particles until a peak level, followed by a shallow exponential decrease until the virus is cleared by the immune system. Mathematically, we model the time of peak viral load as $t_{\textrm{P,p}}=t_{\textrm{T,p}}+\frac{v_{\textrm{P,p}}}{v_{\textrm{I,p}}}$, where $t_{\textrm{T,p}}$ is the time of transmission and $v_{\textrm{I,p}},v_{\textrm{P,p}}>0$ are the viral load increase rate and peak, respectively. The logarithmic viral load $v_{p}$ can be described as
@@
Figure 2: Course of viral load (left) and different, corresponding viral sheds (right) over 20 days. The presented viral sheds are given for varying factor $s_{\textrm{f,p}}$ in (2), ranging from 0.01 (blue) to 0.28 (red).
@@
While an early systematic review [56] and, e.g., [57] found higher SARS-CoV-2 viral loads associated with more severe COVID-19, a more recent systematic review [58] was inconclusive and found contradictory studies on the relationship between viral load and COVID‐19 severity. Accordingly, we model symptoms independently of the viral load. However, the time of symptom onset was observed to be close to the time of peak viral load [59, 60]. When setting parameters, this is taken into account.
@@
Figure 3: Infection state transition model. One-way arrows represent the possible transitions from one infection state to another. Rounded gray borders indicate noninfectious states while red angular borders indicate symptom states that are used for infectious agents only. Orange to violet colors indicate the symptomatic severity and dashed borders indicate if a person is considered as symptomatic (for testing schemes).
@@ 2.5.2 Transmissions
In order to model the transmission of an infection between agents, we consider the viral shed as the proxy of the infectiousness of an agent. This approach is advantageous because viral shedding directly reflects the amount of virus being released by an individual, which has been shown to have a large impact on transmission potential [61]. Unlike symptom-based measures, which may lag behind or be absent in asymptomatic cases, viral shedding captures infectiousness even before symptoms appear or in cases where individuals remain asymptomatic throughout the infection [62, 59]. This proxy also allows for a more granular and dynamic assessment of infectiousness over time. As studies [59, 60] have observed that viral load peaks close to symptom onset, symptom states alone would underestimate the transmission risk during the early, highly infectious phase [55, 63].
@@
where $s_{\textrm{f,p}}$ is the personal viral shed factor for the agent $p$. The viral shed factor is drawn randomly from a probability distribution for each agent and can vary greatly, c.f. [63, Fig. 3]; also see Fig. 2. Similarly to the viral load above, we neglect any viral shed outside of the time interval $[t_{\textrm{T,p}},t_{\textrm{C,p}}]$. Furthermore, we consider only agents which are not Susceptible, Exposed, Recovered and Dead to be infectious. For instance, this means that potential viral loads larger zero are cut on symptomatic recovery. As this leads to a potential cutoff of $s_{\textrm{p}}(t)$, this needs to be taken into account, when determining $\alpha$ and $\beta$. In our simulations, cutting viral loads for recovered people happens mostly when the viral shed is already low, thus leading to a minimal loss of total infectiousness.
@@ 10             Draw new course of infection
Agent A’s exposure rate is further adjusted by seasonality and NPIs, resulting in a final exposure rate of $0.128$ for Agent A. Given that Agent A is wearing a mask that provides $m_{\textrm{p}}=85$ protection ($0.151-m_{\textrm{p}}(t)=$ 0.15), the individual infection rate for Agent A is calculated as $0.038$, as detailed in Table 2. This infection rate will be used to determine the likelihood that Agent A contracts the virus in this specific time step. The implementation of above mentioned NPIs will be explained in detail in the following section.
@@ 2.6.1 Testing and Isolation
Figure 4: An exemplary Testing Strategy with three different Testing Schemes and different Testing Criteria (middle part, blue). We define Testing Schemes for specific locations or, more generally, for different location types (top). Each time an agent enters one of these locations, the Testing Criteria are evaluated (middle). If the criteria are met, the agent undergoes testing with a specified probability and test type. The resulting test remains valid for a predetermined duration. Additionally, the testing strategy can be configured to be active only for a specified time period (bottom).
@@
Figure 5: Example of Testing Schemes 1 and 2 from Fig. 4. In Testing Scheme 1, agents are obliged to test themselves before going to work. It is sufficient to use a Rapid-Antigen Test if they do not show symptoms. All agents who are not symptomatic, i.e., the first three agents in the bottom row, do an antigen test from which the first is false-negative, the second positive and the third negative. Through Testing Scheme 2, a PCR-Test is required if the agent shows symptoms. The last remaining, symptomatic agent in the bottom row performs a PCR test which turns out to be positive. The two positive tests resulted in quarantining at home.
@@ 2.6.2 Venue restrictions and closures
First, the model allows to completely close (particular) locations. Second, it is possible to adjust the capacity of locations, and third, to adjust the likelihood of particular agents to move to a location. By this, we can model a reduced percentage of persons to enter, such as only 50 % of pupils attending school on a particular day. If a capacity threshold is reached, further agents are not allowed to enter the location. For every location, the particular contact matrix can be scaled to simulate contact reduction and social distancing.
@@ 2.6.3 Masks and Compliance
Masks are a crucial NPI to reduce the spread of the virus but their overall effectiveness is highly dependent on correct wearing and compliance [69]. Therefore, our model also implements the usage of different types of Masks such that locations can require agents to use a particular mask – in this case stricter or more effective masks can also be worn to enter a location. In our simulations, a mask reduces the amount of virus particles emitted from an infectious agent and, on the other hand, it reduces the amount of virus particles received by other agents. For details, see (3) and (5). Through this implementation on the individual level, the effectiveness of mask wearing by highly vulnerable persons could be studied with the software. We, furthermore, introduced individual compliance to mask wearing. With the maximum negative compliance value, a person never adheres to any mask mandate, while persons with maximum positive compliance wear a mask even if it is not mandatory. Furthermore, while compliance can be agent-specific, one agent can have different compliance values for wearing masks in different location types.